OpenStudy (anonymous):

uranium hexafluoride is a solid at room temp, but it boils at 56 degrees celsius. determine the density of uranium hexafluoride at 60 degrees celsius and 745 torr.

OpenStudy (anonymous):

P1/T1 = P2/T2 use the equation to find unknown (make sure you match units)

OpenStudy (aaronq):

use PV=nRT you know that n=m/MW and d=m/V substitute into formula PV=mRT/MW ... m/V= d = (PxMW))/RT plug and chug